OpenStudy (datanewb):

What does it mean to raise a vector to the power -2/5? Example below.

OpenStudy (datanewb):

I randomly ended up reading the following thread: http://openstudy.com/users/neemo#/updates/5025dd96e4b07eea166b1db7 The answer and an explanation is given at the bottom. On the surface, I expected to at least understand the problem, as I'm familiar with linear algebra. However, after reading the thread and the answer a couple of times. I still don't understand parts of the question. In particular, what do the following lines mean? \[\left\{ \alpha, \beta, \gamma \right\} = \left\{ 1/6, 1/3\sqrt2,1/2\sqrt3 \right\}\] \[\left\{ \alpha, \beta, \gamma \right\}^{(-2/5)} = 6\] How does one take the exponential power of a vector? As far as I knew, this is not even possible.

OpenStudy (anonymous):

It's not \(\{\alpha,\beta,\gamma\}^{-\frac{2}{5}}\), it's \((\alpha\beta\gamma)^{-\frac{2}{5}}\).

OpenStudy (datanewb):

Oh my, I feel silly now. Thank you.